In the following, find the indicated limit if it exists. Sketch the graph of the function.
f(x) = {4x, x(>or equal to)1/2
{1/x, x<1/2
limf(x)
x =-1/2

Answer by stanbon(75887)

In the following, find the indicated limit if it exists. Sketch the graph of the function.
f(x) = {4x, x(>or equal to)1/2
{1/x, x<1/2
limf(x)
x =-1/2
-----------
Since you are taking the limit as x goes to -1/2, which is < 1/2,
apply the limit to (1/x).
--------------
lim(as x goes to -1/2) of (1/x) = 1/(-1/2) = -2
=========================================================
Cheers,
Stan H.